OpenStudy (anonymous):

COuld someone help me with Geometry B please? The equation of line CD is y = -2x - 2. Write an equation of a line parallel to line CD in slope-intercept form that contains point (4, 5). (4 points) y = -2x + 13 y = x + 7 y = x + 3 y = − 2x − 3

OpenStudy (anonymous):

Hi. I think I can help.

OpenStudy (anonymous):

Okay Thanky you! i'm having a hard time with this i've been thinking its C but i think i'm wrong

OpenStudy (anonymous):

Thank you*

OpenStudy (anonymous):

the slope of parallel lines are equal so m=-2

OpenStudy (texaschic101):

y = -2x - 2 In y = mx + b form, the slope is in the m position. So the slope of this equation is -2. A parallel line will have the same slope. Now we use y = mx + b slope(m) = -2 (4,5)..x = 4 and y = 5 now we sub 5 = -2(4) + b 5 = -8 + b 5 + 8 = b 13 = b so your parallel equation is : y = -2x + 13

OpenStudy (anonymous):

then plug in the point (4,5) to the y=mx+b equation to find the y-intercept, so 5=4(-2) + b..simplify so b=13...thus, your equation is y=-2x + 13

OpenStudy (anonymous):

Okay! Thank you so much! I've been stuck on this one for awhile:/